OpenStudy (anonymous):

Can someone please help me? This is an Algebra 1 question.

OpenStudy (anonymous):

Write five word phrases to be translated using the various key terms from the lesson. Examples must include variables, and there must be at least one example with an addition, subtraction, multiplication, division, parenthesis and exponent, and one with a combination of phrases or symbols.

OpenStudy (anonymous):

5 Keywords are: Commutative Property of Addition or Multiplication, Associative property of Addition or Multiplication, Distributive Property, Reflexive Property of Equality, and Symmetric Property of Equality

OpenStudy (anonymous):

@Broskishelleh can you help me?

OpenStudy (anonymous):

I don't understand what the question means.

OpenStudy (broskishelleh):

@ImaGuest What do you understand?

OpenStudy (anonymous):

I understand the key terms, but I don't know a good way to translate the terms

OpenStudy (broskishelleh):

An example would be for Commutative property of addition: x + 4 Associative property of Addition: a = (b + c) / b = (a + c)

OpenStudy (broskishelleh):

Distributive Property: 2(x + 5) = 2x + 10

OpenStudy (broskishelleh):

Reflexive Property of Equality: a = a/ b = b

OpenStudy (anonymous):

Ohhhhh, I feel stupid now :P

OpenStudy (anonymous):

Thanks

OpenStudy (broskishelleh):

Symmetric Property of Equality: If a = b, b = a

OpenStudy (broskishelleh):

@ImaGuest Understand?

OpenStudy (anonymous):

Yeah, I do.

OpenStudy (anonymous):

Do you think you can translate these to words?

OpenStudy (anonymous):

Sorry it took me a while, I lost connection to the internet because of a storm

OpenStudy (broskishelleh):

I'm sorry @ImaGuest I was watching this: http://www.youtube.com/watch?v=C9Y2voAvVz4. I can help

OpenStudy (anonymous):

I love Mario. I'm getting Mario Cart 8 next week. I can't wait!

OpenStudy (broskishelleh):

Really? Well I like old Mario rather than new, and I like speed runs

OpenStudy (anonymous):

Yeah, the new Mario cart looks good. Personally though, Mario 64 is the best in my opinion.

OpenStudy (broskishelleh):

I guess so, anyways need any more help?

OpenStudy (anonymous):

No, that'll be it for now. Thank you. Do you not Mario 64?